OpenStudy (anonymous):

What is the slope of the line passing through (4,9) and (-1,12)? A) -3/5 B) 7/1 C) 3/5 D) 5/3

OpenStudy (anonymous):

The answer is A!

OpenStudy (anonymous):

Thx

OpenStudy (anonymous):

y2-y1 ------- = 12-9 / -1/4 = -3/5 x2-x1

OpenStudy (anonymous):

Your welcome!